An object initially at rest is subjected to a constant net force. Measurements are taken of its velocity v at different distances d from the starting position. A graph of which of the following should exhibit a straight-line relationship?
The answer is d versus v^2 but i need to know why
Answer #1

the answer is D. Since the force is constant, then acceleration is constant.

v2 = vo2 + 2ad = 0 + 2ad  v2  d


answered by: Mr. Lei